### v1

# Improving Change Captions through Pixel-Level Change Detection Pseudo-Label Learning

Abstract
The existing methods for Remote Sensing Image Change Captioning (RSICC) perform well in simple scenes but exhibit poorer performance in complex scenes. This limitation is primarily attributed to the model’s constrained visual ability to distinguish and locate changes. Acknowledging the inherent correlation between change detection (CD) and RSICC tasks, we believe pixel-level CD is significant for describing the differences between images through language. Regrettably, the current RSICC dataset lacks readily available pixel-level CD labels. To address this deficiency, we leverage a model trained on existing CD datasets to derive CD pseudo-labels. We propose an innovative network with an auxiliary CD branch, supervised by pseudo-labels. Furthermore, a semantic fusion augment (SFA) module is proposed to fuse the feature information extracted by the CD branch, thereby facilitating the nuanced description of changes. Experiments demonstrate that our method achieves state-of-the-art performance and validate that learning pixel-level CD pseudo-labels significantly contributes to change captioning. Our code will be available at: https://github.com/Chen-Yang-Liu/Pix4Cap
Index Terms—  Change captioning, change detection, pseudo-label learning
1 Introduction
Remote sensing image (RSI) change interpretation technology has played a pivotal role in providing valuable insights into environmental changes, urban development, and disaster monitoring [1]. The generation of linguistic information describing changes through RSI change captioning (CC) has emerged as a promising avenue for enhancing the interpretability and communicative capacity of remote sensing imagery.
The existing RSICC methods predominantly adopt the encoder-decoder architecture [2, 3]. Pioneering efforts by Chouaf and Hoxha et al. [4, 5] delved into the RSICC task, employing CNN-based visual encoders for bi-temporal image feature extraction and RNN-based decoders for the sequential generation of words. Extending the RSICC landscape, Liu et al. [6] proposed a large LEVIR-CC dataset and the RSICCformer model, leveraging a dual-branch Transformer to enhance bi-temporal features and highlight the changes of interest. Addressing the multi-scale challenge, PSNet [7] extracts multi-scale features by combining progressive difference perception layers and scale-aware enhancement modules. Chang et al. [8] proposed an attentive encoder comprising hierarchical self-attention blocks and a residual block to extract visual embeddings. PromptCC [9] efficiently utilizes a large language model as a language generator through a multi-prompt learning strategy determined by a coarse change classifier.
Fig. 1: Illustration of our Pix4cap method. Left: the overview of our model. The model comprises two branches: change captioning and auxiliary change detection. The SFA module acts as a pivotal bridge between the two branches. Right: the structure of some important modules.
Despite significant advancements, current RSICC methods face challenges in achieving robust performance, particularly in scenarios involving subtle building changes, challenging lighting conditions, and low-contrast images. The degradation in performance can be largely attributed to the model’s constrained visual capabilities, preventing accurate discernment of pertinent changes while disregarding irrelevant ones. Given the persistence of these challenges, it becomes imperative to explore innovative strategies that can improve RSICC model performance.
Acknowledging the intrinsic connection between change detection (CD) and RSICC, we propose that integrating pixel-level CD outcomes holds substantial promise. Pixel-level CD, involving the identification and localization of changes at the pixel level, provides valuable cues for generating nuanced and accurate descriptions of differences between images. However, a significant hurdle to realizing this potential lies in the absence of pixel-level CD labels within the current RSICC dataset.
To address this crucial gap, we propose to leverage a model pre-trained on existing CD datasets to derive CD pseudo-labels. This pseudo-labeling strategy serves as a stepping stone toward enhancing the model’s capability to perceive changes accurately, particularly in scenes where previous RSICC methods encounter difficulties. In this paper, we propose a novel RSICC network with an auxiliary CD branch supervised by pseudo-labels. Furthermore, we present a Semantic Fusion Augment (SFA) module to seamlessly integrate feature information from the CD branch into the caption branch. Experiments demonstrate that our method achieves state-of-the-art performance and that CD pseudo-label learning is an effective improvement strategy for the RSICC task.
2 Methodology
2.1 Overview
Fig. 1 illustrates the overall structure of our proposed Pix4Cap (Pixel change detection for change Captioning). Pix4Cap comprises two branches dedicated to CD and RSICC, respectively. An SFA module acts as a pivotal bridge, facilitating the integration of features extracted by the CD branch into the RSICC branch and contributing to the generation of change captions. Besides, the CD branch is supervised by the pseudo-labels.
2.2 Change Detection Branch
The architecture of the CD branch mirrors that of UNet. Bi-temporal images serve as inputs to the backbone network, yielding four-scale bi-temporal features $X_{t_{1}}^{i}$, $X_{t_{2}}^{i}$ (i=1,2,3,4). Subsequently, the bi-temporal fusion (BTF) module processes these features alongside their differential features to highlight changes of interest. The CD head then predicts the change map by progressively decoding multi-level features from the BTF module. The structure of the BTF module and CD head are illustrated in Fig. 1 (a) and (b).
2.3 Change Caption Branch
We anticipate that the CD branch will enhance the visual perception capability of the backbone, and the features from the BTF module serving as the neck of the CD branch can provide valuable supplementary information for the change captioning. To this end, we propose an SFA module based on the multi-head attention mechanism [10] to seamlessly integrate the feature $X_{f}^{4}$ from the last BTF module into the RSICC branch, as illustrated in Fig. 1 (c), where $D$ represents the following process:
$X_{dif}=W(X_{t_{2}}-X_{t_{1}})+b+Cos(X_{t_{1}},X_{t_{2}})$ (1)
where $W$ and $b$ denote the projection matrix and bias of a linear layer, and $Cos$ represents the operation of calculating cosine similarity. The subtraction operation and cosine similarity operation are both employed to better capture distinctive features revealing changes of interest.
Following the SFA module, we employ a Transformer decoder [10] with a residual connection as the caption generator. This decoder utilizes visual embeddings to generate accurate and detailed sentences describing changes between images.
3 Experiments
Table 1: Comparisons experiments on the LEVIR-CC dataset, where the bolded results are the best.
[…]
Capt-Rep-Diff [11] 72.90 61.98 53.62 47.41 34.47 65.64 110.57 64.52
Capt-Att [11] 77.64 67.40 59.24 53.15 36.58 69.73 121.22 70.17
Capt-Dual-Att [11] 79.51 70.57 63.23 57.46 36.56 70.69 124.42 72.28
DUDA [11] 81.44 72.22 64.24 57.79 37.15 71.04 124.32 72.58
MCCFormer-S [12] 79.90 70.26 62.68 56.68 36.17 69.46 120.39 70.68
MCCFormer-D [12] 80.42 70.87 62.86 56.38 37.29 70.32 124.44 72.11
RSICCFormer [6] 84.72 76.27 68.87 62.77 39.61 74.12 134.12 77.65
PSNet [7] 83.86 75.13 67.89 62.11 38.80 73.60 132.62 76.78
Baseline 84.87 76.38 69.15 63.25 39.47 74.08 133.96 77.69
Pix4Cap 85.56 77.08 69.79 63.78 39.96 75.12 136.76 78.91
Fig. 2: The comparison between one of five reference captions and the caption generated by our model.
3.1 Experimental Setup
We conduct experiments on the LEVIR-CC dataset [6], a large RSICC dataset with 10077 image pairs and 50385 captions. Among these, 5038 image pairs are changed, while 5039 image pairs are unchanged. Each image pair is annotated by five captions.
Our model is implemented on PyTorch and trained using a single NVIDIA RTX 4090 GPU. The pre-trained BIT model [13] is utilized to generate CD pseudo-labels for each image pair of the LEVIR-CC dataset. ResNet32 [14] is employed as the backbone to extract visual features. The cross-entropy loss function is used to compute both CD loss and caption loss.
[…]
3.2 Experiment Results
Quantitative Results: we conduct ablation experiments and report the quantitative comparison results between our method and previous RSICC methods, including four LSTM-based methods (Capt-Rep-Diff [11], Capt-Att [11], Capt-Dual-Att [11], DUDA [11]) and four Transformer-based methods (MCCFormers-S [12], MCCFormers-D [12], RSICCFormer [6], PSNet [7]). The experimental results are shown in Table 1. Compared with our Pix4Cap, the baseline removes the CD branch and the first cross-attention operation of the SFA module. The ablation result verifies that CD pseudo-label learning can effectively improve the RSICC, resulting in a $1.22\%$ improvement on the $S_{m}^{*}$ metric. Additionally, we can see that our Pix4Cap model achieves SOTA performance.
Quantitative Results: Fig. 2 shows two examples of captions generated by our method. Despite illumination changes and low contrast between buildings and the surrounding environment, our method can accurately distinguish changes of interest and describe the changed object category, their spatial positions, and the relationship between the objects.
4 Conclusion
This paper addresses the significant challenge of weak change recognition in current RSICC techniques, particularly in complex scenes. To overcome this obstacle, we leverage a pre-trained CD model to generate CD pseudo-labels for image pairs in the RSICC dataset. We propose the innovative Pix4Cap model, featuring an auxiliary CD branch supervised by pseudo-labels and a seamlessly integrated SFA module. Our approach ensures the effective integration of pixel-level CD information into the RSICC branch. The experiments validate the effectiveness of learning CD pseudo-labels in improving overall RSICC performance.